# Quantum Universality in Composite Systems: A Trichotomy of Clifford Resources

### v1

@@ VI Case III: Composite Dimensions
We now turn to the final regime: composite dimensions $d=d_{1}d_{2}\cdots d_{n}$ where the factors are pairwise coprime. Unlike prime or prime-power dimensions, which typically require the direct injection of diagonal non-Clifford gates (such as the qubit $T$-gate or the qudit $T_{s}$), composite dimensions offer a path to universality driven by “structural mixing.”
In this regime, the necessary resources are not diagonal non-Clifford gates, but rather standard entangling gates (generalized CNOTs) that couple different coprime factors: intra-qudit gates. While these coupling gates are external to the local Clifford product, they are “natural” resources in any register-based architecture: they are permutation matrices that assist entanglement generation. The central result of this Section is the demonstration that the arithmetic interplay between coprime dimensions, coupled by intra-qudit gates, spontaneously generates the non-Clifford diagonal resources required for universality.
@@ VI.1 Irreducibility via Tensor Mixing
This result generalizes the seminal theorem of Brylinski and Brylinski [Brylinski2002], which established that adjoining any entangling gate to the local unitary group $\mathrm{SU}(d)\otimes\mathrm{SU}(d)$ yields a universal set for two-qudit quantum computation.
@@ VI.3 Main Theorem: Universality
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension decomposed into pairwise coprime prime power factors (i.e., $d_{k}=p_{k}^{m_{k}}$ with $\gcd(d_{i},d_{j})=1$ for $i\neq j$). The group generated by the single-qudit Clifford group and the intra-qudit gates; that is, the group
@@
    $S>\frac{\pi(d_{k}-1)}{2\arcsin(1/4)}.$
    Consider the worst-case scenario with the smallest possible coprime proper powers: $d_{k}=4$ and $d_{j}=9$. The bound for $d_{k}=4$ is $S>6.22(3)\approx 18.66$ and $S=4\times 9=36$. Since $36>18.66$, the condition is strictly satisfied. Any other pair of proper powers yields a larger product, satisfying the bound even more robustly. Therefore, the group is dense in the subsystem $k$.
@@ VI.4 Explicit Universal Gate Sets
A consequence of Theorem 39 is that for composite systems with coprime dimensions, universality is achieved using only standard architectural primitives. In the qubit regime ($d=2$), the group generated by local Clifford gates and CNOTs remains finite (the global Clifford group), requiring the injection of a dedicated magic resource (such as a $T$-gate) to achieve density. In contrast, for coprime composite systems, the inclusion of the intra-qudit gate CNCN—typically considered a standard permutation—is sufficient to break the finite group structure. Due to the arithmetic mismatch between factors, the interplay between the local Clifford groups and the intra-qudit gate generates an infinite, dense group without requiring explicit diagonal magic gates.
@@
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension with pairwise coprime prime power factors. The following finite set of gates constitutes a universal set for quantum computation on arrays of registers of dimension $d$:
@@ VII Discussion and Outlook
The results of this work show that the resources required to supplement the single-qudit Clifford group for universal quantum computation depend fundamentally on the arithmetic structure of the Hilbert space dimension. This dependence is not merely quantitative—involving more or fewer extra gates—but also qualitative: the type of resource changes across the trichotomy. This establishes dimensionality as a computational resource: the factorization properties of the dimension influence not only the cost but the nature of universality.
Our perspective reframes the universality question. Rather than asking which gates generate a dense subgroup of $\mathrm{SU}(d)$, we ask which operations compensate for structural limitations of the Clifford group. Gates that are merely non-Clifford never suffice unless $d=p$, where non-Cliffordness implies magic. Only those that have genuinely “magic” guarantee single-qudit universality. In composite dimensions, these notions diverge: adding a non-Clifford gate can enlarge the group without making it infinite, leaving universality unattained. Achieving universality, therefore, requires gates that both escape the Clifford group and ensure infiniteness. The failure of density may arise from reducibility of the adjoint action, finiteness of the group, or both—and the prime factorization of the dimension dictates which condition applies.
The classification we have presented reveals an unexpected connection between number theory and quantum computation. The prime factorization of the Hilbert space dimension determines whether universality requires explicit magic injection or emerges from structure. This insight highlights dimensionality as a design lever: by choosing coprime or prime-power architectures, one can trade off magic resources against control complexity. This connection is not merely formal; it may have concrete implications for the design of quantum processors.
The broader lesson is that quantum resources are context-dependent. Entanglement, magic, and coherence are necessary for quantum advantage, but which resources are “free” and which must be engineered depends on the computational framework. Dimensionality is context-dependent as well. Its arithmetic structure governs whether magic is intrinsic or must be injected. In homogeneous qubit architectures, entangling gates are free but magic is costly. In coprime architectures, both are free—the price to pay lies in the complexity of controlling heterogeneous subsystems. Understanding these trade-offs across different physical platforms is essential for identifying the most efficient routes to practical quantum computation.
Several limitations and open questions remain. For instance, multi-qudit universality, i.e. density in $\mathrm{SU}(d^{n})$, additionally requires an entangling gate satisfying the Brylinski criterion [Brylinski2002]. For coprime architectures, intra-qudit CNOT gates can serve as both magic and entangling resources, though compilation overhead in heterogeneous registers remains an open question. Also, realizing coprime dimensions physically poses challenges: high-fidelity gates between subsystems of different dimensions, such as a generalized intra-qudit CNOT between a qutrit and a ququart, demand precise control of a twelve-dimensional Hilbert space.
On the other hand, resource-theoretic measures of magic-such as stabilizer nullity [Beverland_2020], mana [Veitch2014], robustness [Howard2017], and stabilizer extent [Bravyi_2019] must be understood when magic emerges from coupling rather than injection. Beyond qudits, whether analogues of our trichotomy exist in continuous-variable systems, which have their own Clifford group and magic resources [Bartlett2002], or in hybrid architectures, is an intriguing question. Additionally, the complexity-theoretic implications of universality in the composite case remain unclear: does adding coprime Clifford + inter-qudit gates interpolate between classically simulable and universal circuits in a meaningful way?
Current architectures typically use homogeneous registers—arrays of identical qubits or qudits—where the Clifford group plus inter-qudit entangling gates remains finite and classically simulable via stabilizer methods [Gottesman1998, AaronsonGottesman2004, Campbell2012], so universality requires magic state distillation [Bravyi2005]. We propose an alternative: heterogeneous coprime architectures, where subsystems of dimensions $d_{1}$ and $d_{2}$ with $\gcd(d_{1},d_{2})=1$ achieve universality using only local Clifford gates and native inter-qudit, coupling. This principle could be realized in trapped ions [RingbauerEtAl2022, Low2020], superconducting circuits [Chicco_2023, Morvan2021], and photonic systems [Wang2020review, liu2024]. The main advantage is eliminating magic state distillation overhead, which dominates resource costs in standard fault-tolerant schemes [Litinski2019, Gidney_2021], even with recent improvements [PrakashSaha2025]. A potential drawback is increased control complexity for heterogeneous registers, but similar challenges already arise in modular and hybrid architectures.
Finally, we note that the trichotomy discovered here may be the first instance of a broader phenomenon. Whenever a mathematical structure admits a notion of “dimension” with nontrivial factorization properties, similar classifications may arise. Exploring such analogies-in continuous-variable systems, in topological quantum computation, or in other algebraic settings-offers an interesting direction for future work.
@@ Acknowledgements.
C.G. was partially supported by Grant INV-2025-213-3452 from the School of Science of Universidad de los Andes. J.R. acknowledges support from the Office of the Vice-president of Research and Creative Activities and the Office of the Faculty of Science Vice-president of Research of Universidad de los Andes under the FAPA grant.

### v2

@@ VI Case III: Composite Dimensions
We now turn to the final case: composite dimensions $d=d_{1}d_{2}\cdots d_{n}$ where the factors are pairwise coprime. Unlike prime or prime-power dimensions, which typically require the direct injection of diagonal non-Clifford gates (such as the qubit $T$-gate or the qudit $T_{s}$), composite dimensions offer a path to universality driven by “structural mixing.”
In this case, the necessary resources are not diagonal non-Clifford gates, but rather standard entangling gates (generalized CNOTs) that couple different coprime factors: intra-qudit gates. While these coupling gates are external to the local Clifford product, they are “natural” resources in any register-based architecture: they are permutation matrices that assist entanglement generation. The central result of this Section is the demonstration that the arithmetic interplay between coprime dimensions, coupled by intra-qudit gates, automatically generates the non-Clifford diagonal resources required for universality.
@@ VI.1 Irreducibility via Tensor Mixing
This result generalizes a theorem by Brylinski and Brylinski [8], which established that adjoining any entangling gate to the local unitary group $\mathrm{SU}(d)\otimes\mathrm{SU}(d)$ yields a universal set for two-qudit quantum computation.
@@ VI.3 Main Theorem: Universality
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension decomposed into pairwise coprime prime-power factors (i.e., $d_{k}=p_{k}^{m_{k}}$ with $\gcd(d_{i},d_{j})=1$ for $i\neq j$). The group generated by the single-qudit Clifford group and the intra-qudit gates; that is, the group
@@
    $S>\pi(d_{k}-1)\big/(2\,\arcsin(1/4)).$
    Consider the worst-case scenario with the smallest possible coprime proper powers: $d_{k}=4$ and $d_{j}=9$. The bound for $d_{k}=4$ is $S>6.22(3)\approx 18.66$ and $S=4\times 9=36$. Since $36>18.66$, the condition is strictly satisfied. Any other pair of proper powers yields a larger product, satisfying the bound strictly. Therefore, the group is dense in the subsystem $k$.
@@ VI.4 Explicit Universal Gate Sets
A consequence of Theorem 48 is that for composite systems with coprime dimensions, universality is achieved using only standard architectural primitives. In the qubit case ($d=2$), the group generated by local Clifford gates and CNOTs remains finite (the global Clifford group), requiring the injection of a dedicated magic resource (such as a $T$-gate) to achieve density. In contrast, for coprime composite systems, the inclusion of the intra-qudit gate CNCN—typically considered a standard permutation—is sufficient to break the finite group structure. Due to the arithmetic mismatch between factors, the interplay between the local Clifford groups and the intra-qudit gate generates an infinite, dense group without requiring explicit diagonal magic gates.
@@
Let $d=d_{1}d_{2}\cdots d_{n}$ be a proper composite dimension with pairwise coprime prime-power factors. The following finite set of gates constitutes a universal set for quantum computation on arrays of registers of dimension $d$:
@@ VII Discussion and Outlook
The results of this paper establish a classification of single-qudit universality, when supplementing the Clifford group, based on the prime factorization of the local Hilbert space dimension $d$. Depending on the value of $d$, a trichotomy arises capturing a hierarchy of algebraic obstructions within the Clifford group: maximality and irreducibility for prime dimensions, reducibility in the case of prime-power dimensions, and automatic resolution via coprime coupling in composite dimensions.
Let us summarize the main findings. The Clifford group $\mathcal{C}_{d}$ is a maximal finite subgroup of $\mathrm{PSU}(p)$ and has irreducible adjoint representation if and only if $d$ is a prime (Theorem 25 and Theorem 24). This maximality and irreducibility imply that any non-Clifford gate—whether diagonal or permutation—suffices for universality (Corollary 27). For prime-power $d=p^{m}$ with $m\geq 2$, the adjoint representation becomes reducible (Theorem 30), fragmenting the Lie algebra into invariant subspaces indexed by divisibility. Universality requires gates that restore irreducibility: either diagonal gates $T_{s}$ with $s\nmid d$ satisfying the bound of Theorem 33, or non-affine permutations such as the transposition (0 1)(0\;1) analyzed in Theorem 39. For coprime composite $d=d_{1}\cdots d_{k}$, the intra-qudit CNOTs connecting subsystems of coprime dimension automatically generate the required non-Clifford resources via Bézout’s identity (Proposition 47).
We emphasize that non-affine permutations constitute a distinct class of universal resources, available only for $d\geq 4$. For $d\in\{2,3\}$, every permutation is affine and, hence, a Clifford gate (Remark 10). This explains why diagonal magic gates remain necessary in qubit and qutrit systems, while for $d\geq 4$ universality can be achieved using only Clifford gates and permutations.
The present work focuses on single-qudit universality, where the Clifford group is associated to the abelian group $\mathbb{Z}_{d}$. From single-qudit universality, general quantum universality follows via a result of Brylinski and Brylinski (Theorem 2). A natural extension concerns multi-qudit Clifford groups associated to $\mathbb{Z}_{d}^{n}$, corresponding to $n$ qudits of dimension $d$. For $d$ prime, the maximality of the $n$-qudit Clifford group was established by Nebe, Rains, and Sloane [34] using $\ell$-adic methods and the classification of finite simple groups. It would be interesting to determine whether the trichotomy identified here extends to the multi-qudit setting for general $d$, and whether the structure of $\mathbb{Z}_{d}^{n}$ for composite $d$ introduces new phenomena beyond those captured by the single-qudit analysis presented here.
From a resource-theoretic perspective, existing measures of magic [2, 38, 22, 6] were developed primarily for qubit systems with explicit $T$-gate injection. The fact that permutation gates—which are classical reversible operations—serve as universal resources for $d\geq 4$ suggests that the relationship between magic and non-Cliffordness may be more subtle in higher-dimensional systems.
On the experimental side, platforms accessing non-binary levels—trapped ions [35, 30], superconducting circuits [10, 33], and photonic systems [41, 28]—provide natural settings for implementing higher-dimensional quantum computation. Our results show that for $d\geq 4$, universality can be achieved without resorting to explicit diagonal magic injection: either via non-affine permutations (for any $d\geq 4$) or through intra-qudit CNOTs (for coprime composite $d$).
A direction for future work is to compare the overhead associated with these high‑dimensional mechanisms against that of traditional magic‑state‑based approaches in low‑dimensional systems, such as magic state distillation [27, 16] or the recently proposed magic state cultivation [17]. Determining whether these alternative routes reduce fault‑tolerant resource costs ultimately hinges on the relative difficulty of implementing and controlling such gates. We leave such an analysis for future research.
@@ Acknowledgements.
The authors thank Vsevolod I. Yashin and Xingshan Cui for their valuable comments and suggestions. C.G. was partially supported by Grant INV-2025-213-3452 from the School of Science of Universidad de los Andes. J.R. acknowledges support from the Office of the Vice-president of Research and Creative Activities and the Office of the Faculty of Science Vice-president of Research of Universidad de los Andes under the FAPA grant.